### v1

@@ 4.6 Overall Architecture of the Decision-Making System
5 Conclusions
The advent of concentrated-liquidity AMMs has transformed liquidity provision from a passive activity into an active portfolio management problem: sustainable returns now hinge on consciously selecting price ranges and reallocating capital in a timely manner. Against this backdrop, the task of optimal liquidity provision (LP) is gaining both practical relevance and academic interest.
This study develops, implements, and validates a general-purpose framework for optimal placement of LP in decentralized exchanges with concentrated liquidity. A key distinguishing feature is a custom backtesting approach that reconstructs historical liquidity solely from executed swaps, approximating the per-range liquidity profile to a preset tolerance and thereby eliminating the need for large on-chain state archives. The resulting workflow is light-weight, easy to deploy, and readily scalable. Beyond strategy design and evaluation, the same tool-set can be used to explore alternative AMM architectures.
The empirical focus lies on fee optimization: historical $\tau$-reset strategies are derived offline, market-condition feature vectors are constructed at each allocation event, and several ML modules are trained and stacked into an integral ensemble. Out-of-time (OOT) experiments on multiple Uniswap V3 pools demonstrate that the ML-driven strategy consistently outperforms a uniform benchmark and that the framework seamlessly transfers between different fee tiers and asset pairs, highlighting its generality and its potential for advancing liquidity management research in the DeFi ecosystem.
External validation against Fractal, an alternative backtesting suite that relies on historical per-range liquidity, shows excellent agreement: the fee estimates differ by less than 4% under identical settings—further supporting the soundness of the proposed methodology.
The broader question of optimizing the live LP position, including hedging and hybrid liquidity placement schemes, has been left for future work. The present framework provides the technical foundation for this endeavor. An open source Python implementation, GPU-accelerated for high-throughput simulations, is available and is intended to foster community adoption and the development of ML-based tooling for DEX optimization.
@@ 5 Conclusions
Appendix Appendix A The state of total liquidity positions versus accumulated fees during the out-of-time (OOT) simulation period
@@
$p_{a_{i}},p_{b_{i}}$ Lower/upper bounds of the $i$-th bucket $B_{i}\in\beta$
@@
$\tau$ General parameter of the $\tau$-reset strategy
@@
$\bm{S}_{e_{i}}^{\Sigma}$ Vector of token-wise pool fees in epoch $e_{i}$
@@
$\bm{W}_{e_{i}}^{\mathrm{LP}}=(w_{1}^{LP},\dots,w_{N}^{LP})$ Bucket-wise vector of $W_{e_{i}}^{\mathrm{LP}}$
@@
$\bm{S}_{e_{i}}^{\mathrm{LP}}$ Vector of token-wise LP fees for epoch $e_{i}$
@@
$\mathbf{S}_{e_{i}}^{\tau}$ Family of random $\tau$-reset strategies
@@
$\bm{\rho}_{e_{i}}^{\mathrm{LP}}$ $(2\tau{+}1)-$size vector of non-zero elements of $\bm{\alpha}_{e_{i}}^{\mathrm{LP}}$
$\bm{\hat{\rho}}_{e_{i}}^{\mathrm{LP}}$ $(\tau{+}1)-$size vector of symmetry-reduced $\bm{\rho}_{e_{i}}^{\mathrm{LP}}$
@@
$\varphi_{e_{i}}^{\text{mOPT}}$ ML-predicted optimal strategy for epoch $e_{i}$
$m$ Number of sub-epochs in epoch $e_{i}$
@@
LRF Liquidity-rule form constraint
PAA Predictive-advantage area

### v2

@@ 5.6 Comparison with an Alternative Tool
[…]
5.7 Results and Analysis
Our experimental evaluation demonstrates the effectiveness of the proposed methodology across multiple dimensions:
Liquidity Approximation Accuracy. Across all tested pools, our parametric reconstruction method achieved historical fee approximation errors of, on average, about 2%, validating the reliability of our approach even without access to historical liquidity data. This represents a significant practical advantage, as complete historical liquidity states are often unavailable or costly to obtain.
Machine Learning Strategy Performance. The integrated ML ensemble consistently outperformed uniform benchmark strategies in the predictive-advantage area (PAA), typically corresponding to $\tau$ values between 5 and 10. For the USDC/ETH 0.3% pool at $\tau=5$, the ML strategy achieved 15.5% higher fees than the uniform benchmark, while at $\tau=10$ the improvement reached 18.4% on the OOT period with given range size parameter $d$ from Table 2. Similar patterns emerged across other pools, with improvements ranging from 13% to 23% depending on the pool characteristics and $\tau$ parameter.
Parameter Sensitivity Analysis. We identified a clear trade-off between liquidity concentration and profitability. While tighter concentration ($\tau=1,2$) often underperformed due to impermanent loss and downside risk, and broader allocation ($\tau\geq 40$) converged toward uniform strategy performance, intermediate values ($\tau=5,10$) provided the optimal balance, maximizing ML model advantage.
Pool Characteristics Impact. More liquid pools (e.g., USDC/ETH 0.05%) showed greater absolute fee improvements and extended PAA ranges, suggesting that our methodology scales effectively with pool activity. The framework maintained consistent performance across diverse asset pairs including stablecoins (USDC/USDT) and volatile pairs (WBTC/ETH).
Modified Strategy Enhancements. By introducing asymmetric $\tau+\eta^{\mathrm{down}}$-reset strategies, we demonstrated further performance improvements. For the pool USDC/ETH 0.05% with $d=10$ USDC, $\tau=5$ and $\eta^{\mathrm{down}}=20$, the modified strategy achieved 88.9% annualized returns compared to 51.9% for buy-and-hold, while reducing maximum drawdown from 16.2% to 11.8%, and significantly improving over the “As-Is” strategy for $\tau=5$ with annualized returns of $92.0\%$.
Validation Against Alternative Tools. Comparison with the Fractal backtesting library showed close agreement (error within 4%), providing external validation of our liquidity approximation methodology despite its independence from historical liquidity data.
These results collectively demonstrate that our framework provides a robust, practical solution for optimizing liquidity provision in CLMMs, with consistent outperformance across diverse market conditions and pool characteristics.
@@ 5.8 Overall Architecture of the Decision-Making System
6 Discussion and Implications
6.1 Theoretical Implications for Market Microstructure
Our findings have several important implications for market microstructure theory in decentralized exchanges.
The consistent outperformance of moderate $\tau$ values over both narrower and broader allocations demonstrates the existence of an optimal concentration level that balances fee capture against impermanent loss risk. This finding parallels inventory management decisions by traditional market makers, who must balance profit from bid-ask spreads against adverse price movement risk. The similarity suggests that despite different institutional arrangements, fundamental economic principles govern liquidity provision decisions across market structures.
The success of our machine learning approach–which uses centralized exchange data to inform decentralized market strategies–highlights the interconnectedness of these markets. This finding contributes to ongoing debates about price discovery in decentralized versus centralized exchanges by demonstrating that informational efficiency in one venue can be leveraged for strategic advantage in another.
The sensitivity of strategy performance to $\tau$ parameters has implications for market stability. If many LPs adopt similar optimal parameters, liquidity could become concentrated in similar price ranges, potentially decreasing the price impact for small or medium sized trades but increasing price impact for significantly large trades. This concentration effect warrants further study regarding market resilience during periods of high volatility. The effect should be taken into account in further studies considering the mutual influence of liquidity providers on each others’ strategies.
6.2 Practical Implications
Our framework provides LPs with actionable tools for strategy development. The identification of predictive-advantage areas shows that the $\tau$-reset strategy can be tuned for specific market conditions, while asymmetric strategy benefits ($\tau$$\eta$ modifications) provide mechanisms for managing downside risk. Practically, LPs can implement these insights through automated strategy execution, potentially improving risk-adjusted returns.
AMM protocol designers can use our findings to inform future developments. The demonstrated importance of balancing concentration benefits against transaction costs and impermanent loss suggests protocols might benefit from mechanisms that help LPs optimize this balance automatically. Additionally, our liquidity reconstruction methodology could be integrated into protocol analytics, improving transparency.
As institutional participation in DeFi grows, our framework offers a rigorous approach to evaluating liquidity provision as an asset class. The ability to quantify historical performance accurately, account for relevant costs, and optimize strategy parameters provides institutional investors with analytical tools needed for informed allocation decisions.
6.3 Limitations and Future Research
While our framework advances the analysis of liquidity provision in CLMMs, several limitations warrant acknowledgment.
Our current model simplifies interactions among multiple LPs. Future work could incorporate game-theoretic elements to analyze strategic interactions more completely, potentially using mean-field approaches or agent-based modeling.
Our analysis focuses on single-protocol strategies. In practice, LPs often deploy capital across multiple protocols simultaneously. Developing frameworks for cross-protocol optimization represents an important extension.
As the boundary between DeFi and traditional finance continues to blur, research exploring integrated strategies–such as using traditional derivatives to hedge impermanent loss–will become increasingly valuable.
Our findings regarding optimal concentration levels could inform the design of next-generation AMM protocols. Research exploring how protocol mechanisms might automatically optimize liquidity distribution based on market conditions could significantly improve capital efficiency and market quality.
7 Conclusion
The transformation of liquidity provision from passive activity to active portfolio management in concentrated liquidity AMMs represents a significant development in market microstructure. This evolution creates both challenges and opportunities for market participants, requiring new analytical tools for performance evaluation and strategy optimization.
This paper has developed and validated a comprehensive framework for addressing these challenges. Our parametric liquidity reconstruction methodology enables accurate historical performance analysis even when complete liquidity data is unavailable, addressing a critical practical limitation in DeFi research. Applying this framework to $\tau$-reset strategies across multiple market conditions, we identify consistent patterns in optimal strategy parameters and demonstrate the potential for machine learning to enhance strategy performance.
Beyond methodological contributions, our findings offer important insights into the economics of automated market making. The existence of predictive-advantage areas in strategy parameter space reveals how liquidity providers can balance competing risks and returns. The effectiveness of asymmetric strategy modifications demonstrates how targeted design choices can mitigate specific risks. And the interconnectedness of centralized and decentralized markets, evidenced by our machine learning approach, highlights the complex informational relationships in modern financial ecosystems.
As decentralized finance continues to evolve and integrate with traditional financial systems, frameworks like the one developed here will become increasingly important for participants, researchers, and regulators alike. By bridging methodological innovation with financial economic insights, this work contributes to our understanding of how market making adapts to new technological paradigms while remaining grounded in fundamental economic principles.
@@ 7 Conclusion
Appendix Appendix A The state of total liquidity positions versus accumulated fees during the out‑of‑time (OOT) simulation period
@@
$p_{a_{i}},p_{b_{i}}$ Lower/upper bounds of the ii‑th bucket $B_{i}\in\beta$
@@
$\tau$ General parameter of the $\tau$‑reset strategy (number of ranges)
@@
$\bm{S}_{e_{i}}^{\Sigma}$ Vector of token‑wise pool fees in epoch $e_{i}$
@@
$\bm{W}_{e_{i}}^{\mathrm{LP}}=(w_{1}^{LP},\dots,w_{N}^{LP})$ Bucket‑wise vector of $W_{e_{i}}^{\mathrm{LP}}$
@@
$\bm{S}_{e_{i}}^{\mathrm{LP}}$ Vector of token‑wise LP fees for epoch $e_{i}$
@@
$\mathbf{S}_{e_{i}}^{\tau}$ Family of random $\tau$‑reset strategies
@@
$\bm{\rho}_{e_{i}}^{\mathrm{LP}}$ $(2\tau{+}1)$-size vector of non‑zero elements of $\bm{\alpha}_{e_{i}}^{\mathrm{LP}}$
$\bm{\hat{\rho}}_{e_{i}}^{\mathrm{LP}}$ $(\tau{+}1)$-size vector of symmetry‑reduced $\bm{\rho}_{e_{i}}^{\mathrm{LP}}$
@@
$\varphi_{e_{i}}^{\text{mOPT}}$ ML‑predicted optimal strategy for epoch $e_{i}$
$m$ Number of sub‑epochs in epoch $e_{i}$
@@
$\eta$ Parameter of the symmetric strategy (number of empty ranges)
$\eta^{\mathrm{up}},\eta^{\mathrm{down}}$ Parameters of the asymmetric strategy (number of empty ranges)
LRF Liquidity‑rule form constraint
PAA Predictive‑advantage area